# Improving governance outcomes through AI documentation: Bridging theory and practice

### v1

@@ 4.2. Potential Impacts & Challenges
Challenges. One barrier to documentation facilitating best practices in development and governance is that the documentation process must be interconnected with other organizational processes to serve as an effective forcing function. In cases where documentation is ad-hoc and relies on tools that are not well-integrated with practitioners’ workflows (Bhat et al., 2023; Heger et al., 2022), connecting documentation processes and artifacts to other aspects of the development and governance process becomes even more challenging. Without a solid documentation infrastructure that aligns with other organizational functions, documentation may fail to effectively enhance development and risk management.
@@
Another challenge is that while high-quality documentation can improve governance and development, poor-quality documentation can have negative impacts. Documentation should prevent redundant efforts, maximize time spent managing unique risks, and support proactive system development. However, these benefits depend on the quality of the documentation. Inadequate or incorrect documentation can lead to wasted time, overlooked risks, and unexpected issues that may require correction later in development. This could result in harms remaining unaddressed or inefficiencies that undermine the overall governance process.
@@ 4.3. Design considerations
How organizations choose to document AI systems reflects both their normative values and practical constraints. Design and implementation choices are also influenced by established and emerging external requirements, such as regulations that companies must adhere to. As organizations (and regulators) aim to maximize the benefits of documentation across various types of AI systems and components, they must often navigate tensions in their approach. Our analysis identified common design tensions or considerations organizations must navigate. These include: balancing customization and standardization, tailoring documentation to specific audiences, determining the appropriate level of detail, deciding on the degree of automation in the documentation process, and incorporating interactivity into documentation artifacts. To an extent, these design considerations are interrelated: the level of detail documentation artifacts contain can also be a form of audience tailoring, for example. Yet each of these design tensions has partially distinct motivations and implications, so we address them separately.
@@
At the organizational level, customization might also be necessary. For example, a healthcare startup may require a different documentation approach than a large, well-established financial institution. In some cases, it may be appropriate to customize documentation at multiple levels to respond to both the organization’s nature and the type of system it is developing (Chmielinski et al., 2020; Gebru et al., 2021; Holland et al., 2018; Richards et al., 2020; Roman et al., 2023; Stoyanovich and Howe, 2019).
@@
A more diffuse but critical benefit of standardization is its role in helping the AI community converge on norms of practice and communication (Arnold et al., 2019; McMillan-Major et al., 2021, 2024). Practitioners often struggle with how to complete documentation or what level of detail to include (Chang and Custis, 2022; Pushkarna et al., 2022; Heger et al., 2022; Miceli et al., 2021), leading them to rely heavily on existing examples, which may not always be appropriate. In one study, practitioners frequently copied documentation from previous projects into their current work, sometimes including irrelevant information or omitting crucial details (Pushkarna et al., 2022). Standardization can mitigate these issues by providing clear expectations for what information should be included and what users should expect to find in documentation artifacts, thus encouraging more consistent practices (McMillan-Major et al., 2021; Pushkarna et al., 2022; Roman et al., 2023).
@@
The choice between customization and standardization is not always a strict binary but a spectrum, allowing for a combination of both. For example, organizations can create standardized templates that are adapted to different contexts, such as healthcare models versus those for software engineers, ensuring consistency while meeting specific needs. However, organizations must thoughtfully decide when to standardize and when to customize.
@@ 4.3.2. Design consideration #2: Determining how much to tailor documentation artifacts to specific audiences
When designing documentation strategies, organizations must consider the diverse needs of multiple stakeholder groups. Different stakeholders often require distinct forms of documentation to be effective. To meet these varied needs, organizations might choose to produce tailored documentation for each stakeholder group. Alternatively, organizations might opt for a single, general-purpose artifact that all stakeholders can use, though this might not be adequately tailored to a specific purpose.
@@
Tailored documentation can also be more actionable for practitioners. One common issue with general-purpose documentation is that it may not clearly guide stakeholders on how to use the information provided (Adkins et al., 2022; Crisan et al., 2022; Stoyanovich and Howe, 2019). This issue is particularly significant when practitioners fail to connect documentation with the ethical impacts of their work (Heger et al., 2022), indicating a need for prescriptive guidance to manage AI risks effectively. The actions practitioners take in response to documentation often depend on their specific roles for two reasons. First, when documentation is aligned with the purpose of their role, practitioners are better able to understand the necessary actions to manage risk. For example, information about how well a system meets user needs may be more actionable for a UX professional than a machine learning engineer. Second, practitioners typically have the authority to act in specific ways within organizations. For instance, compliance teams cannot implement changes to the production codebase, and data scientists are not responsible for conducting legal reviews. Tailored documentation that aligns with each group’s roles and responsibilities can guide them toward actions they are empowered to take.
@@
Furthermore, single-format documentation may be easier to produce and maintain. If practitioners are required to create multiple forms of documentation to meet the needs of different stakeholders, organizations must ensure sufficient time is allocated for this process. Without adequate time, practitioners may rush, leading to low-quality documentation. In cases where time constraints are a concern, it may be more effective to focus on producing a single, high-quality document that is regularly updated. At a minimum, all documentation should clearly indicate its last update to help users gauge its relevance and accuracy.
@@ 4.3.3. Design trade-off #3: Determining the level of interactivity documentation should support
Producing multiple forms of documentation is one way to support a variety of goals, stakeholders, and systems and goals; however, these multiple forms of documentation create challenges for organizations to manage. Interactive documentation offers an alternative solution by enabling stakeholders to access the specific information they need while maintaining overall usability and comprehension.
@@
Empirical research supports the idea that interactive documentation can significantly improve understanding, particularly among non-technical audiences (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023). Tools like Hugging Face’s model inference API444https://huggingface.co/docs/api-inference/ or OpenAI’s developer playground555https://platform.openai.com/playground allow practitioners to engage with models and systems directly, fostering an intuitive understanding of how inputs and outputs are connected. This hands-on interaction is crucial for designing products that leverage the system’s capabilities and for assessing potential risks (Liao et al., 2023). Interactive access is particularly valuable when exploring pre-trained, general-purpose models, whose functionality can vary significantly depending on the implementation (Liao et al., 2023; Moore et al., 2023; Crisan et al., 2022).
@@
Concise documentation can help stakeholders focus on the details that are most relevant to their needs, reducing the likelihood of information overload. For documentation aimed at downstream consumers, including internal risk management professionals, it is essential to provide enough detail to support decision-making without overwhelming them. Information overload can lead to selective attention, where decision-makers may focus on certain details while neglecting others, potentially degrading the quality of their decisions (Phillips-Wren and Adya, 2020), particularly under time constraints (Hahn et al., 1992). For example, faced with extensive and complex documentation on AI systems, consumers might miss critical information or become frustrated and abandon the effort. As one participant in a research study noted, ”I would lose patience after 30-40 seconds if I have to put a lot of effort into finding what I’m looking for” (Crisan et al., 2022). Concise documentation that highlights the most important information for a given stakeholder group can help internal decision-makers manage risk more effectively by ensuring they focus on critical details.
Moreover, concise documentation is often easier to produce and maintain, which can enhance accuracy and encourage more frequent use. Exhaustive documentation may be burdensome to create and keep updated, particularly if the process is manual. For instance, in the initial case studies of an extensive data documentation framework, practitioners reported that completing the documentation took two to three hours, excluding the time needed to gather the required information (Bender and Friedman, 2018). While a few hours might be manageable for a single project, this time commitment becomes overwhelming when scaled across an entire organization, potentially involving hundreds or thousands of datasets, models, or systems. In such scenarios, spreading effort thinly across all systems might detract from the focus on high-risk or high-impact components, undermining critical risk management efforts.
@@
4.4. Recommendations
@@
Given that organizations have unique requirements and operate in vastly different contexts, there is no one-size-fits-all approach to documentation or a universal set of information that applies to all datasets, models, or systems. However, our analysis of proposed frameworks and empirical findings suggests several considerations that organizations should take into account when designing and implementing their documentation strategies. Each of these recommendations could be further strengthened by novel empirical research about how organizations can maximize the benefits of implementing the recommendations in applied contexts.
Identify critical stakeholders and their documentation needs. When developing AI documentation strategies, organizations should begin by identifying the stakeholders who will frequently and significantly interact with both the documentation processes and artifacts. It’s important to consider not only the needs of those who will consume the documentation but also the expertise and capacity of those responsible for producing it. If the documentation is designed to be highly informative for downstream users but is impractical for upstream practitioners to produce, the approach is likely to fail in practice. In such cases, organizations may need to expand their documentation teams to ensure they can meet the needs of consumers effectively. Also, organizations should take into account the technical and non-technical backgrounds of their stakeholders to strike a balance between providing the necessary level of detail and maintaining accessibility.
Evaluate operational resources and constraints and incentivize the production of robust documentation within that context. A realistic assessment of available resources—such as time, personnel, and expertise—is crucial for supporting effective documentation. Overly ambitious documentation efforts without adequate resources can lead to failures, such as incomplete or inaccurate documentation. It is often better to produce concise, high-quality documentation than to attempt a comprehensive approach that falls short. Organizations should ensure that practitioners understand their documentation responsibilities and that incentives are in place to promote thorough and accurate documentation practices.
@@
Leverage empirical evidence and feedback for continuous improvement. Organizations should evaluate potential documentation approaches via pilot projects and periodically collect empirical data on the efficacy of their documentation practices in meeting the needs of relevant stakeholders. Organizations should plan to assess the usability of documentation frameworks and their effectiveness in achieving the organization’s goals (Berman et al., 2024), and adapt their documentation strategies by conducting continuous assessments as their needs or technology evolves. Given the limited empirical data on the effectiveness of documentation approaches, organizations can also contribute to the broader AI community by publicly sharing their findings on what works best in different contexts. Organizations should recognize that empirical evidence and key performance indicators can be qualitative and quantitative. Whereas quantitative evidence can help organizations assess general trends at scale, qualitative evidence can help them understand unique issues in greater depth.
Use standardization as the default approach, and intentionally deviate from it when necessary. Standardization offers many advantages. It allows practitioners to compare documentation artifacts, facilitates search and discoverability of information, establishes consistent expectations for documentation producers and consumers, and can help institutionalize norms of practice. For these reasons, some level of standardization in documentation processes and artifacts is likely to be helpful for most organizations for most AI systems.
@@
To balance standardization and customization, organizations might develop customizable documentation templates or interchangeable modules (Bhat et al., 2023; Heger et al., 2022; Hind et al., 2019; Holland et al., 2018; McMillan-Major et al., 2021; Pushkarna et al., 2022; Richards et al., 2020). Templates could, for example, allow for customization of subsections while maintaining consistent content and structure within those subsections. For instance, companies may include a subsection on the labeling process for AI systems based on supervised learning but exclude it for unsupervised AI systems since they do not require labels. Where companies choose to embrace a greater degree of customization, they should ensure that the benefits of this customization outweigh the benefits of maintaining standardization.
Leverage automation for information extraction and incorporate manual processes for scoped critical synthesis. Some helpful information for documentation, such as dataset composition, statistical properties, model training methods, parameters, and software library or third-party API usage, can be extracted from AI systems using automated procedures, contributing to consistent, high-fidelity documentation of these system elements. While automated documentation can provide reliable and up-to-date information, it has limitations. The aspects of systems that are easiest to document automatically are not always the most critical, and errors in automated documentation can have a greater negative impact than manual errors, as they can persist through each update. Therefore, organizations should be selective about the information included in documentation artifacts, even if it can be easily generated automatically.
@@
Consider interactive documentation interfaces for general-purpose models in particular. Documenting the characteristics of general-purpose models and their potential risks is particularly challenging due to their wide range of possible applications. When downstream practitioners cannot directly communicate with upstream model developers, interpreting the general documentation in the context of specific applications can be difficult. Several studies have shown that interfaces allowing practitioners to observe system outputs in response to given inputs can enhance their understanding of model capabilities and risks (Liao et al., 2023; Crisan et al., 2022; Moore et al., 2023). Consequently, interactive interfaces can serve as a valuable complement to traditional, static documentation.
Focus initially on core information needs for documentation and evaluate potential gaps over time. When designing and implementing new documentation strategies, organizations should start by identifying the essential pieces of information that will meet the needs of key stakeholders. In other words, they should start by designing a concise version of documentation. They should plan to consult with stakeholders who use this documentation to determine whether any additional information is necessary to perform their essential duties and whether any of the initial information provided in documentation artifacts is unnecessary. They should also incorporate learnings from empirical research to augment documentation as needed. Organizations can use this agile approach to manage the information in documentation artifacts, ensuring that the time spent producing and using documentation is maximally efficient and effective at serving stakeholder requirements.
@@
Institutionalize responsible AI and risk management practices beyond documentation. Ethical deliberation is a skill that practitioners can develop over time with intentional practice (Vallor et al., 2018), but organizations need to implement additional interventions beyond documentation to help cultivate this skill. Even when practitioners engage in documentation processes, they may overlook critical risks or fail to recommend appropriate mitigations without a deeper understanding of how AI can harm diverse individuals and communities (Madaio et al., 2024; Chang and Custis, 2022). Therefore, organizations should strive to embed responsible AI and risk management efforts throughout their AI development practices. When responsible AI is integrated into the organization’s culture and broader practices, documentation processes that prompt practitioners to reflect on risks are more likely to be informed by a comprehensive understanding necessary for effective risk management. Conversely, relying solely on documentation to encourage critical reflection is unlikely to be sufficient and may even lead to negative outcomes. Documentation of potential ethical harms will likely be more effective when it focuses on concrete system properties, such as data containing personally identifiable information or existing societal biases, rather than on speculative concerns, which can be subjective and heavily influenced by the expertise and perspectives of the documentation producers.
4.5. Conclusion

### v2

@@ 4.3.3. Challenge #3: Interpersonal and organizational constraints can limit ethical decision making
Beyond undermining the quality of documentation artifacts themselves, poor technical and organizational integration of documentation processes with other organizational infrastructure could limit the extent to which documentation serves as a forcing function for best practices in development and governance generally. In cases where documentation is ad-hoc and relies on tools that are not well-integrated with practitioners’ workflows (Bhat et al., 2023; Heger et al., 2022), connecting documentation processes and artifacts to other aspects of the development and governance process becomes even more challenging (Ahlawat et al., 2024). Without a solid documentation infrastructure that aligns with other organizational functions, documentation may fail to effectively enhance development and risk management.
@@
Another challenge is that while high-quality documentation can have compounding benefits to governance and development, poor-quality documentation can have compounding negative impacts. Documentation should prevent redundant efforts, maximize time spent managing unique risks, and support proactive system development. However, these benefits depend on the quality of the documentation. Inadequate or incorrect documentation can lead to wasted time, overlooked risks, and unexpected issues that may require correction later in development. This could result in harms remaining unaddressed or inefficiencies undermining the overall governance process or quality and reliability of development.
@@ 4.4. Design considerations
How organizations choose to document AI systems reflects both their normative values and practical constraints. Design and implementation choices are also influenced by established and emerging external requirements, such as regulations that companies must adhere to. As organizations (and regulators) aim to maximize the benefits of documentation across various types of AI systems and components, they must often navigate tensions in their approach. Our analysis identified common design tensions or considerations organizations must navigate. These include: balancing customization and standardization, tailoring documentation to specific audiences, determining the appropriate level of detail, deciding on the degree of automation in the documentation process, and incorporating interactivity into documentation artifacts. To an extent, these design considerations are interrelated: the level of detail documentation artifacts contain can also be a form of audience tailoring, for example. And, audience tailoring can be viewed as a form of customization. Yet each of these design tensions at least partially distinct motivations and implications, so we address them separately.
@@
At the organizational level, customization might also be necessary. For example, a healthcare company may require a different documentation approach than a financial company, and the approach either should take may depend on the company’s maturity level (Dotan et al., 2024). In some cases, it may be appropriate to customize documentation at multiple levels to respond to both the organization’s nature and the type of system it is developing (Chmielinski et al., 2020; Gebru et al., 2021; Holland et al., 2018; Richards et al., 2020; Roman et al., 2023; Stoyanovich and Howe, 2019).
@@
A more diffuse but critical benefit of standardization is its role in helping the AI community converge on norms of practice and communication (Arnold et al., 2019; McMillan-Major et al., 2021, 2024). Practitioners often struggle with how to complete documentation or what level of detail to include (Chang and Custis, 2022; Pushkarna et al., 2022; Heger et al., 2022; Miceli et al., 2021), leading them to rely heavily on existing examples, which may not always be appropriate. Standardization can mitigate these issues by providing clear expectations for what information should be included and what consumers should expect to find in documentation artifacts, thus encouraging more consistent practices (McMillan-Major et al., 2021; Pushkarna et al., 2022; Roman et al., 2023).
@@
The choice between customization and standardization is not always a strict binary but a spectrum, allowing for a combination of both. For example, organizations can create standardized templates that are adapted to different contexts, such as healthcare models versus those for software engineers, ensuring consistency while meeting specific needs (Bhat et al., 2023; Heger et al., 2022; Hind et al., 2019; Holland et al., 2018; McMillan-Major et al., 2021; Pushkarna et al., 2022; Richards et al., 2020). However, organizations may face trade-offs when deciding on the level of standardization to employ in documentation.
@@ 4.4.2. Design consideration #2: Determining how much to tailor documentation artifacts to specific audiences
Organizations could tailor their documentation to fit specific application contexts or systems. In addition, tailoring documentation to meet the needs of its intended audience can help ensure it achieves the desired governance outcomes. When designing documentation strategies, organizations must consider the diverse needs of multiple stakeholder groups (Micheli et al., 2023). Different stakeholders often require distinct forms of documentation to be effective. To meet these varied needs, organizations might choose to produce tailored documentation for each stakeholder group. Alternatively, organizations might opt for a single, general-purpose artifact that all stakeholders can use, though this might not be adequately tailored to a specific purpose.
@@
Tailored documentation can also be more actionable for practitioners. One common issue with general-purpose documentation is that it may not clearly guide stakeholders on how to use the information provided (Adkins et al., 2022; Crisan et al., 2022; Stoyanovich and Howe, 2019). This issue is particularly significant when practitioners fail to connect documentation with the ethical impacts of their work (Heger et al., 2022), indicating a need for prescriptive guidance to manage AI risks effectively. The actions practitioners take in response to documentation often depend on their specific roles for two reasons. First, when documentation is aligned with the purpose of their role, practitioners are better able to understand the necessary actions to manage risk. For example, information about how well a system meets user needs may be more actionable for a UX professional than a machine learning engineer (Liang et al., 2024). Second, practitioners typically have the authority to act in specific ways within organizations. For instance, compliance teams cannot implement changes to the production codebase, and data scientists are not responsible for conducting legal reviews. Tailored documentation that aligns with each group’s roles and responsibilities can guide them toward actions they are empowered to take.
@@
Furthermore, single-format documentation may be easier to produce and maintain. If practitioners are required to create multiple forms of documentation to meet the needs of different stakeholders, organizations must ensure sufficient time is allocated for this process. Without adequate time, practitioners may rush, leading to low-quality documentation. In cases where time constraints are a concern, it may be more effective to focus on producing a single, high-quality document that is regularly updated. At a minimum, it may be beneficial if documentation artifacts clearly indicate their last update to help practitioners gauge relevance and accuracy.
@@ 4.4.3. Design trade-off #3: Determining the level of interactivity documentation should support
Producing multiple forms of documentation is one way to support a variety of goals, stakeholders, and systems; however, these multiple forms of documentation create challenges for organizations to manage. Interactive documentation offers an alternative solution by enabling stakeholders to access the specific information they need while maintaining overall usability and comprehension.
@@
Empirical research supports the idea that interactive documentation can significantly improve understanding, particularly among non-technical audiences (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023). Tools like Hugging Face’s model inference API555https://huggingface.co/docs/api-inference/ or OpenAI’s developer playground666https://platform.openai.com/playground allow practitioners to engage with models and systems directly, fostering an intuitive understanding of how inputs and outputs are connected. This hands-on interaction is crucial for designing products that leverage the system’s capabilities and for assessing potential risks (Liao et al., 2023). Interactive access is particularly valuable when exploring pre-trained, general-purpose models, where functionality can vary significantly depending on the implementation (Liao et al., 2023; Moore et al., 2023; Crisan et al., 2022).
@@
Concise documentation could help stakeholders focus on the details that are most relevant to their needs, reducing the likelihood of information overload. For documentation aimed at downstream consumers, including internal risk management professionals, it is essential to provide enough detail to support decision-making without overwhelming them. Information overload can lead to selective attention, where decision-makers may focus on certain details while neglecting others, potentially degrading the quality of their decisions (Phillips-Wren and Adya, 2020), particularly under time constraints (Hahn et al., 1992). For example, faced with extensive and complex documentation on AI systems, consumers might not read documentation at all, miss critical information, or become frustrated and abandon the effort. As one participant in a research study noted, ”I would lose patience after 30-40 seconds if I have to put a lot of effort into finding what I’m looking for” (Crisan et al., 2022). Concise documentation that highlights the most important information for a given stakeholder group can help internal decision-makers manage risk more effectively by ensuring they focus on critical details.
Moreover, concise documentation is likely easier to produce and maintain, which can enhance accuracy and encourage more frequent use. Exhaustive documentation may be burdensome to create and keep updated, particularly if the process is manual. For instance, in the initial case studies of an extensive data documentation framework, practitioners reported that completing the documentation took two to three hours, excluding the time needed to gather the required information (Bender and Friedman, 2018). While a few hours might be manageable for a single project, this time commitment becomes overwhelming when scaled across an entire organization, potentially involving hundreds or thousands of datasets, models, or systems. In such scenarios, spreading effort thinly across all systems might detract from the focus on high-risk or high-impact components, undermining critical risk management efforts.
@@
Providing detailed documentation of the rationale for design and development decisions underscores the subjective nature of these choices, which are often the underlying causes of AI-related harms (Yurrita et al., 2023; Sambasivan et al., 2021). Manually created documentation can empower downstream stakeholders—and ultimately users—to scrutinize and contest the treatment they receive from AI systems (Yurrita et al., 2023). By transparently detailing how normative assumptions and subjective decisions are incorporated into system design and development, organizations can enhance accountability, supporting both internal governance and external stakeholder oversight.
@@
5. Limitations
Our research design and analysis have several limitations. First, we did not directly study the implementation of any specific documentation framework. Consequently, our assessment of the utility of our theories of change, and the challenges associated with them, is based on indirect analysis of existing work rather than direct evaluation.
Another limitation is the broad scope we applied to evaluating documentation. Different domains of documentation, such as documentation for voice data versus art data, documentation for specific system components, such as data versus models, and documentation for different stages of development, such as pretraining data versus model fine-tuning data, may have distinct impacts on governance. Moreover, as with other aspects of governance (Rakova et al., 2021; Winecoff and Watkins, 2022), the role of documentation likely varies across industries and may function differently depending on an organization’s size and level of governance maturity (Ahlawat et al., 2024; Rakova et al., 2021). Our broad approach limits the depth of our analysis in addressing these specific nuances.
Lastly, our work did not explore how multiple forms of transparency could support or improve AI governance. Documentation is one of many tools that can provide transparency into AI systems and support internal and external accountability. In some cases, system- or component-level documentation alone may not sufficiently enable practitioners to understand how systems operate, the risks they pose, or why they might behave in unexpected or potentially harmful ways. In these cases, explainability methods or other forms of transparency may be necessary to provide the insight needed to understand risks or resolve outstanding issues.
6. Directions for future research
@@
Given that organizations have unique requirements and operate in vastly different contexts, there is no one-size-fits-all approach to documentation or a universal set of information that applies to all datasets, models, or systems. Our analysis of proposed frameworks and empirical findings suggests several potential pathways organizations could explore to develop documentation practices tailored to their specific needs and constraints. Further design and empirical research in key areas could address the limitations of our study or expand into other promising directions. Such work could provide organizations across varying maturity levels and domains with more specific, actionable guidance for designing documentation processes that strengthen their AI governance.
Leveraging empirical evidence and feedback for continuous improvement. Our sample of publications contained significantly higher number of proposed documentation frameworks compared to empirical evaluations of their effectiveness, highlighting gaps in understanding how theoretical approaches to documentation can be translated into practical, impactful practices within specific organizational contexts. Notably, most empirical findings on documentation were cross-sectional, offering limited insight into how iterative evaluation and refinement could enhance documentation practices over time within AI organizations.
@@
Future empirical research could fill this gap by identifying how organizations can evaluate potential documentation approaches via pilot projects and periodically collect empirical data on the efficacy of their documentation practices in meeting the needs of relevant stakeholders. Research could also inform how organizations can plan to assess the usability of documentation frameworks and their effectiveness in achieving the organization’s goals (Berman et al., 2024), and adapt their documentation strategies by conducting continuous assessments as their needs or technology evolves. Beyond scholarly research efforts, given the limited empirical data on the effectiveness of documentation approaches, organizations can also contribute to the broader AI community by publicly sharing their findings on what works best in different contexts. Organizations may benefit from empirical evidence and key performance indicators that are both qualitative and quantitative. Quantitative metrics are often best suited to capturing broad patterns across many instances but may lack depth for individual phenomena. Qualitative measures provide ”thick” insights for specific instances (Geertz, 2008) but may have limited generalizability.
Identifying unexpected negative consequences of design choices. Empirical findings from researchers or organizations on how theoretically sound design choices—including those discussed in this work—might lead to unintended negative consequences can offer valuable insights for shaping best practices grounded in real-world experiences. For instance, a study on practitioners might reveal that while interactive interfaces help non-technical users understand information, they could hinder technical practitioners from quickly accessing relevant details. Similarly, research on practitioners or publicly available artifacts might show that practitioners often omit critical information, even when documentation is designed to be concise. Such findings could inform revisions to the frameworks and design considerations discussed in this work.
Informing the degree and design of standardization of artifacts. Our research highlights the advantages of standardization including enabling comparison between documentation artifacts, facilitating search and discoverability of information, establishing consistent expectations for documentation producers and consumers, and institutionalizing norms of practice. For these reasons, some level of standardization in documentation processes and artifacts is likely to be helpful for most organizations for most AI systems. However, standardized formats may overlook or insufficiently address the unique characteristics of specific AI system components or categories (e.g., language versus vision models), or application domains (e.g., social media versus healthcare). Future research could benefit more specific applications by providing insight into the appropriate level of standardization and identifying when and to what extent deviations from standardized formats are advantageous.
@@
One approach to balancing standardization and customization that research has begun to explore is having customizable documentation templates or interchangeable modules (Bhat et al., 2023; Heger et al., 2022; Hind et al., 2019; Holland et al., 2018; McMillan-Major et al., 2021; Pushkarna et al., 2022; Richards et al., 2020). Templates could, for example, allow for customization of subsections while maintaining consistent content and structure within those subsections. For instance, companies may include a subsection on the labeling process for AI systems based on supervised learning but exclude it for unsupervised AI systems since they do not require labels. Future research could explore how templates or other customization approaches can best maintain the usability and institutional benefits of standardization, while allowing for enough customization to address unique risks.
Designing documentation for general-purpose AI systems. Although initiatives such as the EU AI Act require general-purpose model providers to supply downstream deployers with certain technical documentation to support their risk management, research suggests that the properties of upstream or base models may not always be relevant to downstream deployments (Weidinger et al., 2023; Solaiman et al., 2023; Qi et al., 2023). Although approaches for documentation general-purpose models are emerging (Bommasani et al., 2023), disconnects between upstream models and downstream applications pose significant challenges for designing effective documentation recommendations for governing general-purpose systems.
@@
For example, when downstream practitioners cannot directly communicate with upstream model developers, interpreting the general documentation in the context of specific applications can be difficult. Several studies have shown that interfaces allowing practitioners to observe system outputs in response to given inputs can enhance their understanding of model capabilities and risks (Liao et al., 2023; Crisan et al., 2022; Moore et al., 2023). Consequently, interactive interfaces can serve as a valuable complement to traditional, static documentation.
Research that illuminates how traditional, static approaches to documentation can be combined with interactive interfaces and even explainability methods may shed light on how documentation of general-purpose systems can best be communicated throughout the AI development lifecycle. Such work may be especially helpful in designing information architectures for documentation for practitioners without deep AI expertise, since these practitioners often face challenges in understanding complex technical information in model documentation (Crisan et al., 2022).
Disentangling the role of different types of documentation in AI governance. Our analysis reveals that AI documentation can broadly support a range of governance objectives through four general mechanisms. However, different types of documentation—such as data, model, system, or process documentation—either individually or in specific combinations, may serve distinct functions. For example, data documentation provides practitioners with insights into how inputs might influence system behavior. Details about the dataset—such as who is represented and what information is contained within data instances—can alert downstream practitioners to potential biases or risks, including the possibility of unfair outcomes or privacy violations.
Model documentation, on the other hand, can help practitioners evaluate the reliability of system performance or make informed decisions about usage, considering factors like environmental impacts, which can be significant even during model inference (Chien et al., 2023). System and process documentation plays a crucial role in providing operational clarity, enabling downstream stakeholders or organizational decision-makers to assess deployment contexts, overarching system behaviors, performance, risks, and the adequacy of ongoing monitoring processes.
To ensure maximum effectiveness, the form and content of documentation artifacts—as well as the processes involved in creating them—should be responsive to the unique impacts and governance challenges associated with different system components. Future research, particularly empirical studies focused on organizational contexts, could clarify how documenting different system components—individually or in combination—supports distinct or overlapping governance goals. Such analysis could also inform design recommendations for different types of documentation, ensuring they collectively maximize their overall impact on governance.
@@
Institutionalizing responsible AI and risk management practices beyond documentation. Ethical deliberation is a skill that practitioners can develop over time with intentional practice (Vallor et al., 2018), but organizations need to implement additional interventions beyond documentation to help cultivate this skill. Even when practitioners engage in documentation processes, they may overlook critical risks or fail to recommend appropriate mitigations without a deeper understanding of how AI can harm diverse individuals and communities (Madaio et al., 2024; Chang and Custis, 2022). Therefore, organizations should strive to embed responsible AI and risk management efforts throughout their tools and AI development practices. When responsible AI is integrated into the organization’s culture and broader practices, documentation processes that prompt practitioners to reflect on risks are more likely to be informed by a comprehensive understanding necessary for effective risk management.
An open research question is how different tools and methods for raising awareness of ethical issues, fostering ethical deliberation, and empowering practitioners to act on these insights can work together to enhance AI governance. For instance, fairness toolkits can help practitioners less-experienced in algorithmic fairness better recognize and deliberate on trade-offs between system fairness and other performance indicators (Balayn et al., 2023). However, the tools and processes organizations adopt to institutionalize risk management practices may inadvertently narrow practitioners’ focus, limiting their ability to address emergent problems that fall outside the scope of these tools and processes or encouraging them to apply risk management techniques in inappropriate contexts (Balayn et al., 2023; Ahlawat et al., 2024).
Future research could explore how documentation and other AI risk mitigation tools can enable practitioners to manage risks consistently and rigorously while remaining flexible enough to identify and address new and unforeseen concerns. Research on how explainability methods can support documentation may be especially helpful in this regard. Stakeholders seek explanations for a variety of reasons, such as translating system predictions into concrete actions, ensuring compliance with relevant policies and regulations, understanding the impact of different model factors on their outputs, among others (Suresh et al., 2021). While documentation could alert practitioners to possible issues that might arise in deployment, it does not necessarily enable causal understanding of issues. Explainability methods, in contrast, can support practitioners in understanding model behavior at the instance level, helping them debug models or systems (Bhatt et al., 2020). Explainability methods may be especially useful in for helping practitioners conceptualize the behavior of general-purpose AI systems such as large language models (LLMs) (Liao and Vaughan, 2024). By identifying which stakeholders would benefit most from explanations and what their goals are, organizations can design explainability approaches that work in tandem with documentation to support responsible risk management and decision making.
7. Conclusion